OpenStudy (anonymous):

please explain fan and medal...just how to solve it helps...I like to think for myself Perpendicular lines AB and CD intersect at point E. If m∠AED = x + 20, what is the value of x?

OpenStudy (anonymous):

give me best response first

OpenStudy (anonymous):

are the medal given to Olympians before the competition...hehem no..but if I must

OpenStudy (anonymous):

70 degrees

OpenStudy (anonymous):

The lines are perpendicular. so they intersect at right angles = 90 degrees

OpenStudy (anonymous):

x+20 =90 x=90-20 x= 70 deg. option D

OpenStudy (anonymous):

thank you very much...god bless your day